Thanks chords of contact of the pair of tangents to the circle x^2+y^2=1 draw from any point on the line 2x+y=4 pass through (a,b) , then find a^2+b^2

Grade:11

2 Answers

Arun
6 years ago
Hello Ajay


Equation of chord of contact:
hx+ky=1

(h, k) lie on the line, we have
2h+k=4
hx+(4-2h)y=4

=>(1/2, 1/4)
hence a = ½
b = ¼
now a^2 + b^2 = 1/4  + 1/16 = 5/16
 
Regards
Arun (askIITians forum expert)
Ajay Pachankar
6 years ago
I can`t understand these equation and how solve thisHow we get value of a and b from this plz help me and show how solve properly